# A look at the operator product expansion in critical dynamics

### v1

@@ VII Extension to $O\left(N\right)$ models
$S$ $=$ $\int_{t,x}\bar{\phi}_{0}^{m}\left(t,x\right)\Biggr{\{}\partial_{t}\phi_{0}^{m}\left(t,x\right)+D_{0}\left(-\partial^{2}+m_{0}^{2}+\frac{g_{0}}{3!}\phi_{0}^{n}\left(t,x\right)\phi_{0}^{n}\left(t,x\right)\right)\phi_{0}^{m}\left(t,x\right)$
@@ VII.1 OPE at the Gaussian level
$\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{T}$$\equiv$ $\left[\frac{\phi^{a}\phi^{b}}{2}-\frac{\delta^{ab}}{N}\frac{\phi^{m}\phi^{m}}{2}\right]$
$\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{S}$$\equiv$ $\left[\frac{\delta^{ab}}{N}\frac{\phi^{m}\phi^{m}}{2}\right]\,,$
@@
$\phi^{a}\times\phi^{b}$ $=$ $G_{ab}+2\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{T}+2\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{S}+\cdots\,.$
@@ VII.2.1 The composite operator $\left[\frac{\phi^{m}\phi^{m}}{2}\right]$
$c_{\phi\phi\left[\frac{\phi^{m}\phi^{m}}{2}\right]}$$\approx$ $\frac{1}{N}\Biggr{\{}2+g\frac{N+2}{3}\left(\frac{\log\left(\pi x^{2}\mu^{2}\right)+\gamma}{16\pi^{2}}+\frac{1}{\left(4\pi\right)^{2}}\Gamma\left(0,\frac{x^{2}}{4Dt}\right)\right)\Biggr{\}}\,,$ (21)
@@
$c_{\phi_{1}\phi_{1}S}\left(y,x\right)$ $=$ $\frac{1}{x^{2\Delta_{\phi_{1}}-\Delta_{S}}}\sqrt{\frac{2}{N}}\left(1-\frac{N+2}{2\left(N+8\right)}\epsilon+\frac{N+2}{2\left(N+8\right)}\Gamma\left(0,\frac{1}{y}\right)\epsilon\right)+O\left(\epsilon^{2}\right)$
$c_{\phi_{1}\bar{\phi}_{1}S}\left(y,x\right)$ $=$ $-\frac{1}{x^{\Delta_{\phi_{1}}+\Delta_{\bar{\phi}_{1}}-\Delta_{S}}}\frac{\left(N+2\right)}{\sqrt{2N}\left(N+8\right)}\frac{e^{1-\frac{1}{y}}\theta\left(y\right)}{y}\epsilon+O\left(\epsilon^{2}\right)\,.$
The static limit of $c_{\phi_{1}\phi_{1}S}$ matches with the result found in Dey et al. (2017) via the analytic conformal bootstrap.
@@ VIII.1 Action in the large N limit
In this section we rewrite the action of the model A in a way that allows one to take the large $N$ limit straightforwardly. In order to do so, we translate the SDE from Ito to the Stratonovich interpretation and consider the associated path integral. The reason for this is that certain formal manipulations are based on the standard rules of calculus, which are not always valid for the path integral. This is particularly relevant when there are “ordering ambiguities”, i.e., terms such as $\hat{p}\hat{q}$ in quantum mechanics. As we have seen in section IV.1, the Fokker-Planck Hamiltonian contains such terms. Ordering ambiguities are in one-to-one correspondence with the choice of path integral discretization Berezin (1980) and we shall stick with the midpoint prescription, which is associated with Weyl ordering. As reviewed in Appendix B, by using this discretization an integral over a time derivative satisfies the usual Stokes theorem.
To study the large $N$ limit we follow the approach explained in Moshe and Zinn-Justin (2003). The partition function takes the form
$Z$ $=$ $\int{\cal D}\phi{\cal D}\bar{\phi}\exp\Biggr{[}-\int_{tx}\bar{\phi}^{a}\left(\partial_{t}\phi^{a}+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!N}\phi^{m}\phi^{m}\right)\phi^{a}\right)+\int_{tx}D_{0}\bar{\phi}^{a}\bar{\phi}^{a}\Biggr{]}$
×$\times\exp\Biggr{[}\frac{1}{2}D\frac{\delta^{2}H}{\delta\phi^{a}\delta\phi^{a}}\Biggr{]}\,,$
$\exp\Biggr{[}\frac{1}{2}D\frac{\delta^{2}H}{\delta\phi^{a}\delta\phi^{a}}\Biggr{]}$ $=$ $\exp\Biggr{[}\frac{1}{2}D\int_{t,xy}\left(N\left(-\partial_{x}^{2}+r\right)+\frac{u_{0}}{3!}\left(\phi^{2}+\frac{2}{N}\phi^{2}\right)\right)\delta\left(x-y\right)\Biggr{]}\,.$
$Z$ $=$ $\int{\cal D}\phi{\cal D}\bar{\phi}\exp\Biggr{[}-\int_{tx}\bar{\phi}^{a}\left(\partial_{t}\phi^{a}+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!N}\phi^{m}\phi^{m}\right)\phi^{a}\right)+\int_{tx}D_{0}\bar{\phi}^{a}\bar{\phi}^{a}\Biggr{]}$
$S$ $=$ $\int_{tx}\bar{\sigma}\left(\partial_{t}\sigma+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!N}\rho\right)\sigma\right)$
$S$ $=$ $N\Biggr{\{}\int_{tx}\bar{\sigma}\left(\partial_{t}\sigma+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!}\rho\right)\sigma\right)$
@@
00 0 $=$ $D_{0}\left(r_{0}+\frac{u_{0}}{3!}\rho\right)\sigma-2D_{0}\bar{\sigma}$
00 0 $=$ $D_{0}\left(r_{0}+\frac{u_{0}}{3!}\rho\right)\bar{\sigma}+i\lambda\sigma$
00 0 $=$ $-i\frac{1}{2}\left(\rho-\sigma^{2}\right)+\frac{1}{2}\int_{\omega q}\frac{i}{\frac{\omega^{2}+D_{0}^{2}\left(q^{2}+r_{0}+\frac{u_{0}}{3!}\rho\right)^{2}}{2D_{0}}+i\lambda}$
00 0 $=$ $\bar{\sigma}D_{0}\frac{u_{0}}{3!}\sigma-i\frac{\lambda}{2}-\frac{1}{2}D_{0}\left(\int_{q}\right)\frac{u_{0}}{3!}$
@@
00 0 $=$ $\left(r_{0}+\frac{u_{0}}{3!}\rho\right)\sigma$
00 0 $=$ $-\frac{1}{2}\left(\rho-\sigma^{2}\right)+\frac{1}{2}\int_{q}\frac{1}{\left(q^{2}+r_{0}+\frac{u_{0}}{3!}\rho\right)}\,.$
@@
00 0 $=$ $m^{2}\sigma$
00 0 $=$ $\frac{\sigma^{2}}{2}-\frac{3}{u}\left(m^{2}-r_{0}\right)+\frac{1}{2}\int_{q}\frac{1}{\left(q^{2}+m^{2}\right)}\,.$
@@
00 0 $=$ $\frac{3!}{u_{0}}\left(r_{0}-m^{2}\right)+\frac{\Gamma\left(1-d/2\right)}{\left(4\pi\right)^{d/2}}\left(m^{2}\right)^{\frac{d-2}{2}}\,.$
@@
$m^{2}$$\approx$ $\left(-\frac{\left(4\pi\right)^{d/2}}{\Gamma\left(1-d/2\right)}\frac{3!}{u_{0}}r_{0}\right)^{\frac{2}{d-2}}\,\equiv\,m_{{\rm o.s.}}^{2}\,.$
@@ VIII.2 OPE coeffients
$\langle\sigma\sigma\rangle$$\approx$ $\frac{1}{N}\frac{1}{\left(4\pi\right)^{d/2}}\Biggr{\{}2^{d-2}\Gamma\left(\frac{d}{2}-1\right)\frac{1}{x^{d-2}}+\Gamma\left(1-\frac{d}{2}\right)m_{{\rm o.s.}}^{d-2}-2^{d-4}\Gamma\left(\frac{d}{2}-2\right)x^{4-d}m_{{\rm o.s.}}^{2}$
$-\frac{2^{d-2}}{x^{d-2}}\Gamma\left(\frac{d}{2}-1,\frac{x^{2}}{4Dt}\right)+\frac{2^{d-4}}{x^{d-2}}m_{{\rm o.s.}}^{2}x^{2}\Gamma\left(\frac{d}{2}-2,\frac{x^{2}}{4Dt}\right)\Biggr{\}}\,.$
@@
$c_{\phi_{1}\phi_{1}\phi_{2}}$ $=$ $\frac{1}{x^{2\Delta_{1}-\Delta_{2}}}\sqrt{\frac{2}{N}}\sqrt{\frac{\Gamma\left(d-2\right)\sin\left(\pi\frac{d}{2}\right)}{\left(\frac{d}{2}-2\right)\pi}}\frac{1}{\Gamma\left(\frac{d}{2}-1\right)}\left[1-\frac{\Gamma\left(\frac{d}{2}-2,\frac{1}{y}\right)}{\Gamma\left(\frac{d}{2}-2\right)}\right]\,.$
@@
$c_{\phi_{1}\bar{\phi}_{1}\phi_{2}}$ $=$ $-\frac{2}{\sqrt{N}\sqrt{\pi}}\frac{\theta\left(y\right)}{x^{d-2}}e^{1-\frac{1}{y}}y^{1-\frac{d}{2}}\sqrt{\frac{\sin\left(\frac{\pi d}{2}\right)\Gamma\left(d-2\right)}{\left(d-4\right)\Gamma\left(\frac{d}{2}-2\right)^{2}}}\,.$
@@ IX Summary and conclusions
In this work we considered the OPE in the context of critical dynamics and computed some OPE coefficients for the so-called model A. Moreover, we showed how it is possible to compute certain universal quantities that are related to the OPE coefficients by some suitable normalizations. The role of the fluctuation-dissipation theorem for the calculated coefficients has also been discussed.
More precisely, in section IV we derived the Gaussian form of the OPE. We found it instructive to derive it by introducing a normal ordering prescription stemming from interpreting the Fokker-Planck equation as a functional “Schroedinger” equation. The corrections to the Gaussian OPE have been calculated in section V by considering a mass expansion of the correlation function of interest. By defining suitable normalizations we have derived a universal form of the OPE coefficients in section VI. The $O\left(N\right)$ dynamical model has been considered in sections VII and VIII where the perturbation theory and the large $N$ limit have been considered.
It is possible to extend the present study in several directions. On the one hand, it is possible to push our calculations to higher orders and to consider further OPE coefficients. On the other hand, the present paper considers the critical dynamics associated with systems that have already relaxed to equilibrium. Therefore, it would be interesting to generalize our study also to the case of non-thermal fixed points Berges (2015) or to consider the effect of the initial state in more detail Täuber (2014).
@@ Appendix A Conventions
$\int_{\omega}\equiv\int\frac{d\omega}{2\pi}$ and $\int_{q}\equiv\int\frac{d^{d}q}{\left(2\pi\right)^{d}}\,.$
@@
$f\left(t,x\right)$ $=$ $\int_{\omega q}e^{-i\omega t+iqx}\tilde{f}\left(\omega,q\right)$
$\tilde{f}\left(\omega,q\right)$ $=$ $\int_{tx}e^{+i\omega t-iqx}f\left(t,x\right)\,.$
@@
$\langle\phi\left(t,x\right)\phi\left(0,0\right)\rangle$ $=$ $\int_{\omega q}e^{-i\omega t+iqx}\frac{2D}{\omega^{2}+D^{2}\left(q^{2}+m^{2}\right)^{2}}$ (23)
@@
$-2^{d-2}\Gamma\left(\frac{d}{2}-1,\frac{x^{2}}{4Dt}\right)+2^{d-4}m^{2}x^{2}\Gamma\left(\frac{d}{2}-2,\frac{x^{2}}{4Dt}\right)\Biggr{\}}+o\left(m^{2}\right),$
@@
Appendix B Fokker-Planck Hamiltonian, ordering, and mid-point prescription

### v2

@@ VII Extension to $O\left(N\right)$ models
$S$ $=$ $\int_{t,x}\bar{\phi}_{0}^{m}\left(t,x\right)\Biggr{\{}\partial_{t}\phi_{0}^{m}\left(t,x\right)+D_{0}\left(-\partial^{2}+m_{0}^{2}+\frac{g_{0}}{3!}\phi_{0}^{n}\left(t,x\right)\phi_{0}^{n}\left(t,x\right)\right)\phi_{0}^{m}\left(t,x\right)$ (69)
@@ VII.1 OPE at the Gaussian level
$\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{T}$$\equiv$ $\left[\frac{\phi^{a}\phi^{b}}{2}-\frac{\delta^{ab}}{N}\frac{\phi^{m}\phi^{m}}{2}\right]$ (71)
$\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{S}$$\equiv$ $\left[\frac{\delta^{ab}}{N}\frac{\phi^{m}\phi^{m}}{2}\right]\,,$ (72)
@@
$\phi^{a}\times\phi^{b}$ $=$ $G_{ab}+2\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{T}+2\left[\frac{\phi^{a}\phi^{b}}{2}\right]_{S}+\cdots\,.$ (73)
@@ VII.2.1 The composite operator $\left[\frac{\phi^{m}\phi^{m}}{2}\right]$
$C_{\phi\phi\left[\frac{\phi^{m}\phi^{m}}{2}\right]}$$\approx$ $\frac{1}{N}\Biggr{\{}2+g\frac{N+2}{3}\left(\frac{\log\left(\pi x^{2}\mu^{2}\right)+\gamma}{16\pi^{2}}+\frac{1}{\left(4\pi\right)^{2}}\Gamma\left(0,\frac{x^{2}}{4D|t|}\right)\right)\Biggr{\}}\,,$ (74)
@@
$C_{\phi_{1}\phi_{1}S}\left(y,x\right)$ $=\frac{1}{x^{2\Delta_{\phi_{1}}-\Delta_{S}}}\sqrt{\frac{2}{N}}\left(1-\frac{N+2}{2\left(N+8\right)}\epsilon+\frac{N+2}{2\left(N+8\right)}\Gamma\left(0,\frac{1}{|y|}\right)\epsilon\right)$ (76a)
$C_{\phi_{1}\bar{\phi}_{1}S}\left(y,x\right)$ $=-\frac{1}{x^{\Delta_{\phi_{1}}+\Delta_{\bar{\phi}_{1}}-\Delta_{S}}}\frac{\left(N+2\right)}{\sqrt{2N}\left(N+8\right)}\frac{e^{1-\frac{1}{y}}\theta\left(y\right)}{y}\epsilon$ (76b)
up to order $O\left(\epsilon^{2}\right)$. The static limit of $C_{\phi_{1}\phi_{1}S}$ matches with the result found in Dey et al. (2017) via the analytic conformal bootstrap.
@@ VIII.1 Action in the large N limit
In this section we rewrite the action of the model A in a way that allows one to take the large $N$ limit straightforwardly. In order to do so, we translate the SDE from Ito to the Stratonovich interpretation and consider the associated path integral. The reason for this is that certain formal manipulations are based on the standard rules of calculus, which are not always valid for the path integral. This is particularly relevant when there are “ordering ambiguities”, i.e., terms such as $\hat{p}\hat{q}$ in quantum mechanics. As we have seen in section IV.1, the Fokker-Planck Hamiltonian contains such terms. Ordering ambiguities are in one-to-one correspondence with the choice of path integral discretization Berezin (1980) and we shall stick with the midpoint prescription, which is associated with Weyl ordering. As reviewed in Appendix C, by using this discretization an integral over a time derivative satisfies the usual Stokes theorem.
To study the large $N$ limit we follow the approach explained in Moshe and Zinn-Justin (2003). In Appendix D we show that the large $N$ limit is captured by the following action
$S$ $=$ $N\Biggr{\{}\int_{tx}\bar{\sigma}\left(\partial_{t}\sigma+D_{0}\left(-\partial^{2}+r_{0}+\frac{u_{0}}{3!}\rho\right)\sigma\right)$ (77)
@@
$+\frac{1}{2}\mbox{Tr}\log\left(\frac{1}{2D_{0}}\left(\partial_{t}+D_{0}\left(\cdots\right)\right)\left(-\partial_{t}+D_{0}\left(\cdots\right)\right)+i\lambda\right)\Biggr{\}}\,.$
@@
00 0 $=$ $D_{0}\left(r_{0}+\frac{u_{0}}{3!}\rho\right)\sigma-2D_{0}\bar{\sigma}$ (78)
00 0 $=$ $D_{0}\left(r_{0}+\frac{u_{0}}{3!}\rho\right)\bar{\sigma}+i\lambda\sigma$ (79)
00 0 $=$ $-i\frac{1}{2}\left(\rho-\sigma^{2}\right)+\frac{1}{2}\int_{\omega q}\frac{i}{\frac{\omega^{2}+D_{0}^{2}\left(q^{2}+r_{0}+\frac{u_{0}}{3!}\rho\right)^{2}}{2D_{0}}+i\lambda}$ (80)
00 0 $=$ $\bar{\sigma}D_{0}\frac{u_{0}}{3!}\sigma-i\frac{\lambda}{2}-\frac{1}{2}D_{0}\left(\int_{q}\right)\frac{u_{0}}{3!}$ (82)
@@
00 0 $=$ $\left(r_{0}+\frac{u_{0}}{3!}\rho\right)\sigma$ (83)
00 0 $=$ $-\frac{1}{2}\left(\rho-\sigma^{2}\right)+\frac{1}{2}\int_{q}\frac{1}{\left(q^{2}+r_{0}+\frac{u_{0}}{3!}\rho\right)}\,.$ (84)
@@
00 0 $=$ $m^{2}\sigma$ (85)
00 0 $=$ $\frac{\sigma^{2}}{2}-\frac{3}{u}\left(m^{2}-r_{0}\right)+\frac{1}{2}\int_{q}\frac{1}{\left(q^{2}+m^{2}\right)}\,.$ (86)
@@
00 0 $=$ $\frac{3!}{u_{0}}\left(r_{0}-m^{2}\right)+\frac{\Gamma\left(1-d/2\right)}{\left(4\pi\right)^{d/2}}\left(m^{2}\right)^{\frac{d-2}{2}}\,.$ (87)
@@
$m^{2}$$\approx$ $\left(-\frac{\left(4\pi\right)^{d/2}}{\Gamma\left(1-d/2\right)}\frac{3!}{u_{0}}r_{0}\right)^{\frac{2}{d-2}}\,\equiv\,m_{{\rm o.s.}}^{2}\,.$ (88)
@@ VIII.2 OPE coeffients
$\langle\sigma\sigma\rangle$$\approx$ $\frac{1}{N}\frac{1}{\left(4\pi\right)^{d/2}}\Biggr{\{}2^{d-2}\Gamma\left(\frac{d}{2}-1\right)\frac{1}{x^{d-2}}+\Gamma\left(1-\frac{d}{2}\right)m_{{\rm o.s.}}^{d-2}-2^{d-4}\Gamma\left(\frac{d}{2}-2\right)x^{4-d}m_{{\rm o.s.}}^{2}$ (89)
$-\frac{2^{d-2}}{x^{d-2}}\Gamma\left(\frac{d}{2}-1,\frac{x^{2}}{4D|t|}\right)+\frac{2^{d-4}}{x^{d-2}}m_{{\rm o.s.}}^{2}x^{2}\Gamma\left(\frac{d}{2}-2,\frac{x^{2}}{4D|t|}\right)\Biggr{\}}\,.$
@@
$C_{\phi_{1}\phi_{1}\phi_{2}}=\frac{1}{x^{2\Delta_{1}-\Delta_{2}}}\sqrt{\frac{2}{N}}\sqrt{\frac{\Gamma\left(d-2\right)\sin\left(\pi\frac{d}{2}\right)}{\left(\frac{d}{2}-2\right)\pi}}\frac{1}{\Gamma\left(\frac{d}{2}-1\right)}\left[1-\frac{\Gamma\left(\frac{d}{2}-2,\frac{1}{|y|}\right)}{\Gamma\left(\frac{d}{2}-2\right)}\right]\,.$ (90)
@@
$C_{\phi_{1}\bar{\phi}_{1}\phi_{2}}=-\frac{2}{\sqrt{N}\sqrt{\pi}}\frac{\theta\left(y\right)}{x^{d-2}}e^{1-\frac{1}{y}}y^{1-\frac{d}{2}}\sqrt{\frac{\sin\left(\frac{\pi d}{2}\right)\Gamma\left(d-2\right)}{\left(d-4\right)\Gamma\left(\frac{d}{2}-2\right)^{2}}}$ (91)
@@ IX Summary and conclusions
In this work we considered the OPE in the context of critical dynamics and computed some OPE coefficients for the so-called model A. More precisely, we have calculated the leading OPE coefficients of the two-point function and the response function in the case where the fluctuation-dissipation theorem holds. Moreover, we showed how it is possible to compute certain universal quantities that are related to the OPE coefficients by some suitable normalizations. We calculated explicitly such universal quantities in several cases and approximations. These quantities and our discussion of the role of the fluctuation-dissipation theorem for the calculated coefficients constitute the main results of the present work. Since the OPE in the dynamical case has been far less explored than its static counterpart, our work constitutes an exploration of the use of the OPE framework in critical dynamics.
More precisely, in section IV we derived the Gaussian form of the OPE. We found it instructive to derive it by introducing a normal ordering prescription stemming from interpreting the Fokker-Planck equation as a functional “Schroedinger” equation. Within the $\epsilon$-expansion, the leading corrections to the Gaussian OPE have been calculated in section V by considering a mass expansion of the correlation function of interest. By defining suitable normalizations we have derived a universal form of the OPE coefficients in section VI. The $O\left(N\right)$ dynamical model has been considered in sections VII and VIII where the perturbation theory and the large $N$ limit have been considered.
It is possible to extend the present study in several directions. On the one hand, it is possible to push our calculations to higher order in the coupling expansion and to consider further OPE coefficients. On the other hand, it would be interesting to extend our analysis to other dynamical models, such as those considered in the Hohenberg-Halperin classification Hohenberg and Halperin (1977). We do not see any particular obstacle preventing this generalization and we expect that certain aspects, such as the application of FDT, are fully analogous. A natural questions to be addressed here is if any particular constraint applies to the OPE of conserved modes. Furthermore, the present paper considers the critical dynamics associated with systems that have already relaxed to equilibrium. Therefore, it would be interesting to generalize our study also to the case of non-thermal fixed points Berges (2015) or to consider the effect of the initial state in more detail Täuber (2014). It would also be very interesting to understand if there is any method that allows one to determine the OPE coefficients “self-consistently”, i.e., by imposing a number of suitable requirements. This is, however, beyond the scope of the present paper. Finally, the reader may wonder if any of the present results are relevant for experiments. Let us note that the used approximations, i.e., first order in the $\epsilon$-expansion and leading order in the $1/N$-expansion, are too rough to be of any experimental interest. However, if a sufficiently precise result for the coefficients and for the expectation values of the composite operator was obtained, one could in principle compare the results with experiments or simulations by looking at the two-point functions slightly away from the critical temperature or at the four-point function in a suitable momentum configuration.
@@ Appendix A Conventions
$\int_{\omega}\equiv\int\frac{d\omega}{2\pi}$ and $\int_{q}\equiv\int\frac{d^{d}q}{\left(2\pi\right)^{d}}\,.$ (93)
@@
$f\left(t,x\right)$ $=$ $\int_{\omega q}e^{-i\omega t+iqx}\tilde{f}\left(\omega,q\right)$ (94)
$\tilde{f}\left(\omega,q\right)$ $=$ $\int_{tx}e^{+i\omega t-iqx}f\left(t,x\right)\,.$ (95)
@@
$\langle\phi\left(t,x\right)\phi\left(0,0\right)\rangle$ $=$ $\int_{\omega q}e^{-i\omega t+iqx}\frac{2D}{\omega^{2}+D^{2}\left(q^{2}+m^{2}\right)^{2}}$
@@
$-2^{d-2}\Gamma\left(\frac{d}{2}-1,\frac{x^{2}}{4D|t|}\right)+2^{d-4}m^{2}x^{2}\Gamma\left(\frac{d}{2}-2,\frac{x^{2}}{4D|t|}\right)\Biggr{\}}+o\left(m^{2}\right),$
@@
Appendix B Details regarding the composite operator $\left[\phi^{2}/2\right]$
In this section we provide details regarding the calculations and definitions associated with $\left[\phi^{2}/2\right]$.
$Z_{20}$ $=$ $\frac{1}{16\pi^{2}\epsilon}\,.$ (98)
Figure 1: The diagram 1(a) is the diagram contributing to the VEV of $\left[\phi^{2}/2\right]$ at Gaussian order. The diagrams 1(b) contributes at order $O(g)$ while diagram 1(c) vanishes. We do not display here the diagrams associated with the counterterms stemming out of the renormalization factors $Z_{22}$, $Z_{20}$, and $Z_{m^{2}}$.
$Z_{ij}$ $=$ $\sum_{k}Z_{ij,k}g^{k}\,.$ (99)
At the Gaussian level, one finds $Z_{22,0}=1$ and $Z_{20,0}=\frac{1}{16\pi^{2}\epsilon}$. To order $O\left(g\right)$ one obtains the following result for the counterterms
$Z_{22,1}$ $=$ $\frac{1}{\left(4\pi\right)^{2}}\frac{1}{\epsilon}$ (100)
$Z_{20,1}$ $=$ $\frac{1}{256\pi^{4}\epsilon^{2}}\,.$ (101)
The counterterm $Z_{22,1}$ implies the anomalous dimension $\gamma_{\phi^{2}}=-Z_{\phi^{2}}^{-1}\dot{Z}_{\phi^{2}}\approx\frac{g}{16\pi^{2}}$. By using the fixed point value $g_{*}\approx 16\pi^{2}\epsilon/3$, one obtains the scaling dimension $\Delta_{2}$ given in (10).
Finally, we consider the static limit of the two-point of $\left[\phi^{2}/2\right]$ to $O\left(g\right)$. At the fixed point, the diagrams contributing to this order are shown in Figure 3.
After an explicit computation and performing the $\epsilon$-expansion to the first order at the fixed point, one obtains:
${\cal N}_{2}$$\approx$ $4\sqrt{2}\pi^{2}-\epsilon\frac{2}{3}\sqrt{2}\left(2\gamma\pi^{2}-\pi^{2}+2\pi^{2}\log\pi-2\pi^{2}\log\mu\right)\,.$ (102)
Appendix C Fokker-Planck Hamiltonian, ordering, and mid-point prescription